# Continuous-Time Value Iteration for Multi-Agent Reinforcement Learning

### v2

@@ 3.2.2 Policy Learning
While analytical optimal control laws can be derived in some cases by maximizing the Hamiltonian, such closed-form solutions might not be available in complex multi-agent systems such as MPE or MuJoCo. To overcome this challenge, we use an actor network to generate control inputs, replacing the need for analytical controls in the critic network used to approximate value functions. The actor and critic networks are refined iteratively until convergence, enabling the actor network to approximate optimal control policies. During training, we compute a continuous-time advantage function derived by residual $\mathcal{R}_{\theta}(x,t)$. This advantage function is used for policy gradient, where each agent’s decentralized policy $\pi_{\phi_{i}}(u_{i}\mid x,t)$ is optimized to maximize long-term return.
@@
$Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl(x(t+\delta t),t+\delta t\bigr).$ (6)
@@
$A(x,t,u)=-\rho V(x,t)+\partial V_{t}(x,t)+\nabla_{x}V(x,t)^{\!\top}\,f(x,u)+r(x,u).$ (7)
The proof is attached at Appendix A.2.
With the critic’s instantaneous advantage:
$A_{\theta}(x,t,u)=-\rho V_{\theta}(x,t)+\partial_{t}V_{\theta}(x,t)+\nabla_{x}V_{\theta}(x,t)^{\!\top}f(x,u)+r(x,u),$ (8)
we update each agent’s policy network $\pi_{\phi_{i}}(u_{i}\mid x,t)$ in a decentralized fashion. For agent $i$, we minimize the negative expected advantage under the joint policy:
$\mathcal{L}_{p_{i}}=-A_{\theta}\bigl(x,t,u\bigr)\log\pi_{\phi_{i}}\bigl(u_{i}\mid x,t\bigr),$ (9)
@@
$Q^{\pi_{\text{new}}}(x,t,u)\geq Q^{\pi_{\text{old}}}(x,t,u).$ (10)
The proof can be found at Appendix A.3.
@@ 3.3 Value Gradient Iteration Module
The performance of continuous-time control policies depends critically on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x,t)$. Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs [51, 27, 10]. However, the gradients obtained from standard PINN training are often noisy or misaligned with true trajectory behavior. Although prior work [51] improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings. To address this, we introduce the VGI module that iteratively refines the value gradient approximations during training.
@@
$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\!\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$ (11)
The VGI target in Eq. (11) can be interpreted as a one-step unrolling of the Bellman equation in the space of gradients. The first term captures the instantaneous contribution of the local reward gradient, while the second term propagates the downstream value information through the Jacobian of the system dynamics. This construction resembles a semi-discretized version of the value gradient flow and provides a practical surrogate for supervised gradient learning in the absence of ground-truth derivatives. The derivation process is posted at Appendix A.4.
@@
$G(\zeta)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\top}\zeta,$ (12)
and assume the dynamics $\|\nabla_{x}f(x_{t},u_{t})\|$ is bounded . Then $G$ is a contraction, and the sequence $\zeta^{(k+1)}=G(\zeta^{(k)})$ converges to a unique fixed point $\zeta^{*}\in\mathbb{R}^{d}$.
The proof is detailed at Appendix A.5
Rather than introducing a separate network to predict value gradients, we directly compute the automatic derivative of the shared PINN critic $V_{\theta}(x,t)$. This gradient is then trained to match the VGI-generated target defined in Eq. (11). Specifically, we minimize the mean squared error between the computed and target gradients:
$\mathcal{L}_{\text{vgi}}=\bigl\|\nabla_{x}V_{\theta}(x_{t},t)-\hat{g}_{t}\bigr\|_{2}^{2},$ (13)
where $\hat{g}_{t}=\nabla_{x}r_{\phi}(x_{t},u_{t},\Delta t)\Delta t+e^{-\rho\Delta t}\nabla_{x}f_{\psi}(x_{t},u_{t},\Delta t)^{\!\top}\nabla_{x}V_{\theta}(x_{t+\Delta t},t+\Delta t).$ Here, $r_{\phi}(x,u,\Delta t)$ denotes a reward models and $f_{\psi}(x,u,\Delta t)$ represents a dynamics model, where $\phi$ and $\psi$ are respective network parameters.
@@ 3.4.1 Dynamics Model and Reward Model
In a continuous‑time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice. Instead, we adopt a discrete‑time model‑based approach [36, 20, 13] that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
$\mathcal{L}_{\text{dyn}}=\bigl\|f_{\psi}(x_{t},u_{t},\Delta t)-x_{t+\Delta t}\bigr\|_{2}^{2}.$ (14)
After learning $f_{\psi}$, we recover the continuous‑time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
Similarly, we fit a reward network $r_{\phi}(x_{t},u_{t},\Delta t)$ to the observed instantaneous reward $r_{t}$:
$\mathcal{L}_{\text{rew}}=\bigl\|r_{\phi}(x_{t},u_{t},\Delta t)-r_{t}\bigr\|_{2}^{2}.$ (15)
@@ 3.4.2 Anchor Loss for Critic Network
In addition to the HJB residual and terminal-condition losses, we incorporate a TD-style anchor loss to improve both the stability and accuracy of value learning. While the residual loss enforces the correctness of the value gradient, it does not constrain the value of $V(x,t)$. Terminal-condition losses can provide such supervision, but they often rely on access to well-defined terminal targets, which may be unavailable in complex continuous control environments such as MuJoCo. In these cases, the anchor loss offers an additional source of value landscape, helping the critic produce reasonable value approximations even when terminal rewards are sparse, delayed, or difficult to specify. We define the one-step return as:
$R_{t}=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V_{\theta}\bigl(x_{t+\Delta t},\,t+\Delta t\bigr).$ (16)
The anchor loss then enforces the value network to match these returns:
$\mathcal{L}_{\text{anchor}}=\bigl\|V_{\theta}(x_{t},t)-R_{t}\bigr\|_{2}^{2}.$ (17)
The overall critic objective combines all four losses:
$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{res}}}_{\text{HJB residual}}\;+\;\underbrace{\lambda_{\text{term}}\,\mathcal{L}_{\text{term}}}_{\text{terminal BC}}\;+\;\underbrace{\lambda_{\text{anchor}}\,\mathcal{L}_{\text{anchor}}}_{\text{TD anchor}}\;+\;\underbrace{\lambda_{g}\,\mathcal{L}_{\text{vgi}}}_{\text{VGI consistency}}.$ (18)
Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value‐bootstrap anchoring, and gradient consistency. In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory. The detailed training process is listed in Appendix A.6.
@@ 4 Experimental Results
We evaluate our Value Iteration via PINN (VIP) method on two continuous‑time multi‑agent benchmarks: MPE [25] and multi‑agent MuJoCo [37, 31]. In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations. This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP. Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous‑time RL baselines in these environments? (2) VGI ablation: How much does the VGI module contribute to final performance and training stability? (3) PINN design choice: How does activation function choice and loss term weighting in critic network affect the performance of VIP? (4) Time discretization impact: How well do discrete-time and continuous-time methods perform under arbitrary or unfixed time intervals?
@@
Benchmarks. We evaluate our method against baselines across eight experimental settings. We extend the MPE framework to a continuous-time formulation by using a variable-step Euler integration scheme, where the time interval $\Delta t$ is sampled from a predefined range at each step. Experiments are conducted on the cooperative navigation and predator prey environments. Similarly, we adapt MuJoCo to continuous-time settings and evaluate on a suite of multi-agent locomotion tasks, including ant (2 $\times$ 4, 2 $\times$ 4d, 4 $\times$ 2), walker, swimmer, and cheetah. Further implementation details of these benchmarks are provided in Appendix B.2 and B.3. Lastly, we introduce a simple yet illustrative coupled oscillator environment to highlight the behavior of exact value functions, value gradients, and the relative performance of different methods under a controlled setting.
@@ 4.1 Baseline Methods
To evaluate our CT‑MARL framework VIP, we compare against three continuous‑time policy iteration baselines and include an ablated variant of our method without VGI: CT‑MBRL (ODE) [48]: A continuous‑time model‑based RL approach that learns system dynamics via Bayesian neural ODEs. This method uses an actor‑critic framework to approximate state‑value functions and derive the continuous-time optimal policies.
@@
Differential Policy Iteration (DPI) [23]: A model‑based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations. Integral Policy Iteration (IPI) [23]: A partially model‑free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement. In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions. Ablation (w/o VGI): An ablated version of VIP without VGI, which isolates the efficacy of value gradient refinement.
@@ 4.2 Results Analysis
Model Performance. We evaluate all RL methods in MPE and MuJoCo environments using five random seeds and report the mean cumulative reward curves in Fig. 2 and 4. The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL. As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations [22, 51].
@@
To address this limitation, we incorporate the anchor and VGI loss terms in Eq. (18), which capture the landscape of value and its gradient so that PINNs converge to the true values. Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments. This observation is consistent with the conclusion in the previous studies [50], which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
To further demonstrate the importance of VGI, we revisit the didactic example, generate 400 rollouts from sampled initial states, and compute the average value using models with and without VGI. As shown in Fig. 3, the value contour projected onto the $d_{1}$$d_{2}$ frame reveals that the model with VGI closely matches the ground truth in both structure and scale, while the model without VGI produces significantly biased approximations. These results confirm that adding VGI is necessary to improve the accuracy of value approximations. A detailed comparison of the corresponding value gradients is also illustrated in Fig. 3. Furthermore, we evaluate the sensitivity of each loss term for the critic network by measuring the minimum distance to prey in Fig. 6, which highlights the critical role of PINN in refining value and policy networks.
@@
Choice of PINN Design. We first examine the impact of activation function choice when incorporating PINN-based losses into critic network refinement. Specifically, we train VIP using ReLU and Tanh activations in both MuJoCo and MPE environments and report the accumulated rewards in Fig. 5. The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks. This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies [1, 51]. The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs. PINNs typically use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs. The PINN residuals, including PDEs and boundary conditions, are further optimized using gradient descent. Smooth activation functions like Tanh support stable and accurate gradient flow throughout training, enabling more effective value approximations.
@@
In contrast, VIP with ReLU often encounters zero-gradient regions during backpropagation, which results in gradient explosion for deeper network architectures or degrades the learning of value functions due to insufficient nonlinearity. Therefore, the choice of a smooth activation function like Tanh is better suited for physics-informed learning, thereby ensuring more accurate approximations of the value functions.
We also investigate the impact of weight parameters for PINN loss terms during VIP training. Specifically, we evaluate three configurations in Eq. (18): 1) balanced weights for all loss terms; 2) a large weight for boundary condition while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms. As shown in Fig. 7, the best accumulated reward performance is achieved when all loss terms are properly balanced. Imbalanced weight settings yield stiffness in the training dynamics of PINNs [42], which makes the dominant loss term (with the largest weights) converge faster than the others [43].
@@
In our experiments, such an imbalance causes VIP only to satisfy the PDEs or boundary conditions during training, ultimately leading to poor value approximations. Lastly, we evaluate the robustness of VIP and MADDPG [24] by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts. Fig. 8 illustrates that VIP maintains a nearly constant return across different time intervals, whereas MADDPG’s performance degrades significantly as the interval increases. This result highlights the advantage of VIP in continuous-time multi-agent scenarios.
5 Conclusion and Limitations
We propose a novel approach that integrates PINNs into the actor-critic framework to solve CT-MARL problems. Specifically, we approximate value functions using HJB-based PINNs and introduce VGI to improve value approximations, thus mitigating the adverse impact of inaccurate value approximations on policy learning. We validate our VIP across continuous‑time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA. Furthermore, we investigate the importance of activation function choice and loss term weighting for VIP performance. In summary, our proposed VIP offers a promising approach to solve CT-MARL problems. However, we notice the limitation of VIP: the current work only focuses on cooperative CT-MARL tasks. Many real-world multi-agent scenarios involve competitive interactions, which have not yet been explored within CT-MARL frameworks. In competitive settings, agents no longer share a common value function, and such values often become discontinuous due to collision penalties in the reward function, which poses convergence challenges for PINN [51]. As part of future work, we aim to extend VIP to explore competitive CT-MARL problems, particularly those involving state constraints.

### v3

@@ 3.2.2 Policy Learning
While analytical optimal control laws can be derived in some cases by maximizing the Hamiltonian (Nakamura-Zimmerer et al., 2021; Shilova et al., 2024), such closed-form solutions are not available in complex multi-agent systems like MPE or MuJoCo. To overcome this challenge, we use an actor network to generate control inputs, replacing the need for analytical controls in the critic network used to approximate value functions. The actor and critic networks are refined iteratively until convergence, enabling the actor network to approximate optimal control policies. During training, we compute a continuous-time advantage function derived by residual $\mathcal{R}_{\theta}(x_{t})$. This advantage function is used for policy gradient, where each agent’s decentralized policy $\pi_{\phi_{i}}(u_{i}~|~x_{t})$ is optimized to maximize long-term return.
@@
$Q(x_{t},u_{t})=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}V\bigl(x_{t+\delta t}\bigr).$ (5)
@@
$A(x_{t},u_{t})=-\rho V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}\,f(x_{t},u_{t})+r(x_{t},u_{t}).$ (6)
The proof is attached in Appendix A.2.
With the critic’s instantaneous advantage
$A_{\theta}(x_{t},u_{t})=-\rho V_{\theta}(x_{t})+\nabla_{x_{t}}V_{\theta}(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t}),$ (7)
we update each agent’s policy network $\pi_{\phi_{i}}(u_{i}~|~x_{t})$ in a decentralized fashion. For agent $i$, we minimize the negative expected advantage under the joint policy
$\mathcal{L}_{p_{i}}=-A_{\theta}\bigl(x_{t},u_{t}\bigr)\log\pi_{\phi_{i}}\bigl(u_{i}~|~x_{t}\bigr),$ (8)
@@
Since our actor is trained by using the advantage function, it is important to ensure that this update direction leads to policy improvement. The following lemma formalizes this property and shows that a gradient step on our actor loss yields a value-increasing policy update.
@@
$Q^{\pi_{\text{new}}}(x_{t},u_{t})\geq Q^{\pi_{\text{old}}}(x_{t},u_{t}).$ (9)
The proof can be found in Appendix A.3.
@@ 3.3 Value Gradient Iteration Module
The performance of continuous-time control policies depends explicitly on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x)$. Recent studies have demonstrated that the accuracy of the value directly affects the learned control policies (Zhang et al., 2024b). However, enforcing the HJB PDE through residual loss alone does not guarantee accurate value gradients. Because in high-dimensional multi-agent systems, small gradient errors can be significantly amplified by coupled dynamics, leading to inaccurate policy gradients. To address this limitation, we introduce the VGI module, which explicitly propagates value gradients along sampled trajectories and refines the gradient estimates in a self-consistent manner. By iteratively updating $\nabla_{x}V(x)$ using the local dynamics and the current value approximation, VGI provides a trajectory-aligned correction signal that complements the global PDE constraint, resulting in substantially more accurate value-gradient learning.
@@
$\nabla_{x_{t}}V(x_{t})=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f(x_{t},u_{t})^{\!\top}\nabla_{x_{t+\Delta t}}V(x_{t+\Delta t}).$ (10)
The VGI target in Eq. 10 can be interpreted as a one-step unrolling of the Bellman equations in the space of gradients. The first term captures the instantaneous contribution of the local reward gradient, while the second term propagates the downstream value information through the Jacobian of the system dynamics. This construction resembles a semi-discretized version of the value gradient flow and provides a practical surrogate for supervised gradient learning in the absence of ground-truth derivatives. The derivation process is posted at Appendix A.4. To justify why the VGI refinement is stable and converges to a promising gradient estimate, we analyze its update rule as a fixed-point iteration.
@@
$G(\zeta)=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f(x_{t},u_{t})^{\top}\zeta,$ (11)
and assume the dynamics $\|\nabla_{x_{t}}f(x_{t},u_{t})\|$ is bounded . Then $G$ is a contraction, and the sequence $\zeta^{(k+1)}=G(\zeta^{(k)})$ converges to a unique fixed point $\zeta^{*}\in\mathbb{R}^{d}$.
The proof is detailed in Appendix A.5
Rather than introducing a separate network to predict value gradients, we directly compute the automatic derivative of the shared PINN critic $V_{\theta}(x_{t})$. This gradient is then trained to match the VGI-generated target defined in Eq. 10. Specifically, we minimize the mean squared error between the computed and target gradients:
$\mathcal{L}_{\text{vgi}}=\bigl\|\nabla_{x_{t}}V_{\theta}(x_{t})-\hat{g}_{t}\bigr\|_{2}^{2},$ (12)
where $\hat{g}_{t}=\nabla_{x_{t}}r_{\phi}(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f_{\psi}(x_{t},u_{t})^{\!\top}\nabla_{x_{t+\Delta t}}V_{\theta}(x_{t+\Delta t}).$ Here, $r_{\phi}(x_{t},u_{t})$ denotes a reward model and $f_{\psi}(x_{t},u_{t})$ represents a dynamics model, where $\phi$ and $\psi$ are respective network parameters.
@@ 3.4.1 Dynamics Model and Reward Model
In a continuous‑time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice. Instead, we adopt a discrete‑time model‑based approach (Sutton, 1991; Hafner et al., 2019) that we train a neural network $f_{\psi}(x_{t},u_{t})$ to predict the next state $x_{t+\Delta t}$ via
$\mathcal{L}_{\text{dyn}}=\bigl\|f_{\psi}(x_{t},u_{t})-x_{t+\Delta t}\bigr\|_{2}^{2}.$ (13)
After learning $f_{\psi}$, we recover the continuous‑time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t})-x_{t}}{\Delta t}$.
Similarly, we fit a reward network $r_{\phi}(x_{t},u_{t})$ to the observed instantaneous reward $r_{t}$:
$\mathcal{L}_{\text{rew}}=\bigl\|r_{\phi}(x_{t},u_{t})-r_{t}\bigr\|_{2}^{2}.$ (14)
@@ 3.4.2 Anchor Loss for Critic Network
In addition to the HJB residual loss, we incorporate a TD-style anchor loss to improve both the stability and accuracy of value learning. While the residual loss enforces the correctness of the value gradient, it does not constrain the value of $V(x)$. Terminal-condition losses can provide such supervision, but they often rely on access to well-defined terminal targets, which may be unavailable in complex continuous control environments such as MuJoCo. In these cases, the anchor loss offers an additional source of value landscape, helping the critic produce reasonable value approximations even when terminal rewards are sparse, delayed, or difficult to specify. We define the one-step return as
$R_{t}=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V_{\theta}\bigl(x_{t+\Delta t}\bigr).$ (15)
The anchor loss then enforces the value network to match these returns
$\mathcal{L}_{\text{anchor}}=\bigl\|V_{\theta}(x_{t})-R_{t}\bigr\|_{2}^{2}.$ (16)
The overall critic objective combines all four losses
$\mathcal{L}_{\text{total}}=\underbrace{\mathcal{L}_{\text{res}}}_{\text{HJB residual}}\;+\;\underbrace{\lambda_{\text{anchor}}\,\mathcal{L}_{\text{anchor}}}_{\text{TD anchor}}\;+\;\underbrace{\lambda_{g}\,\mathcal{L}_{\text{vgi}}}_{\text{VGI consistency}}.$ (17)
Here $\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE residuals, value‐bootstrap anchoring, and gradient consistency. In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory. The detailed training process is listed in Appendix A.6.
@@ 4 Experimental Results
We evaluate our Value Iteration via PINN (VIP) method on two continuous‑time multi‑agent benchmarks: MPE (Lowe et al., 2017b) and multi‑agent MuJoCo (Peng et al., 2021). In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations. This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP. Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperform existing continuous‑time RL baselines in these environments? (2) VGI ablation: How much does the VGI module contribute to final performance and training stability? (3) PINN design choice: How does activation function choice and loss term weighting in the critic network affect the performance of VIP? (4) Time discretization impact: How well do discrete-time and continuous-time methods perform under arbitrary or unfixed time intervals? (5) Comparison between VIP and discrete-time baselines: Could discrete-time baselines perform comparable at continuous-time settings?
@@
Benchmarks. We evaluate our method against competitive baselines across eight experimental settings involving up to six agents and 113 state dimensions. We extend the MPE framework to a continuous-time formulation by using a variable-step Euler integration scheme, where the time interval $\Delta t$ is sampled from a predefined range at each step. Experiments are conducted on the cooperative navigation and predator prey environments. Similarly, we adapt MuJoCo to continuous-time settings and evaluate on a suite of multi-agent locomotion tasks, including ant (2 $\times$ 4, 2 $\times$ 4d, 4 $\times$ 2), walker, swimmer, and cheetah (6 $\times$ 1). Further implementation details of these benchmarks are provided in Appendix B.2 and B.3. Lastly, we introduce a simple yet illustrative coupled oscillator environment to highlight the behavior of exact value functions, value gradients, and the relative performance of different methods under a controlled setting.
@@ 4.1 Baseline Methods
To evaluate our CT‑MARL framework VIP, we compare against four continuous‑time policy iteration baselines and include an ablated variant of our method without VGI: CT‑MBRL (ODE) (Yildiz et al., 2021): a continuous‑time model‑based RL approach that learns system dynamics via Bayesian neural ODEs. This method uses an actor‑critic framework to approximate state‑value functions and derive the continuous-time optimal policies.
@@
Differential Policy Iteration (DPI) (Lee and Sutton, 2021): a model‑based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations. Integral Policy Iteration (IPI) (Lee and Sutton, 2021): a partially model‑free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement. HJBPPO (Mukherjee and Liu, ): a recent method that employs a PINN-based critic to approximate the HJB residual and leverages a standard PPO-style policy optimization scheme to guide agent learning. In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions. Ablation (w/o VGI): an ablated version of VIP without VGI, which isolates the efficacy of value gradient refinement. In addition, to evaluate how VIP compares with DTRL methods, we incorporate three widely used baselines in our experiments: MATD3 (Zhan et al., 2021), MAPPO (Yu et al., 2022), and MADDPG (Lowe et al., 2017a).
@@ 4.2 Results Analysis
Model Performance. We evaluate all RL methods in MPE and MuJoCo environments using five random seeds and report the mean cumulative reward curves in Fig. 2 and 4. The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL. As traditional time-dependent HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations (Krishnapriyan et al., 2021). This limitation becomes even more severe in the infinite-horizon setting, where the HJB formulation is time-independent and terminal losses are no longer available.
@@
To address this limitation, we incorporate the anchor and VGI loss terms in Eq. 17, which capture the landscape of value and its gradient so that PINNs converge to the true values. Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments. This observation is consistent with the conclusion in the previous studies (Zhang et al., 2024a), which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement. To further demonstrate the importance of VGI, we revisit the didactic example, generate 400 rollouts from sampled initial states, and compute the average value using models with and without VGI.
@@
As shown in Fig. 3, the value contour projected onto the $d_{1}$$d_{2}$ frame reveals that the model with VGI closely matches the ground truth in both structure and scale, while the model without VGI produces significantly biased approximations. These results confirm that adding VGI is necessary to improve the accuracy of value approximations. A detailed comparison of the corresponding value gradients is also illustrated in Fig. 3. Furthermore, we evaluate the sensitivity of each loss term for the critic network by measuring the minimum distance to prey in Fig. 6, which highlights the critical role of PINN in refining value and policy networks.
Choice of PINN Design. We first examine the impact of activation function choice when incorporating PINN-based losses into critic network refinement. Specifically, we train VIP using ReLU and Tanh activations in both MuJoCo and MPE environments and report the accumulated rewards in Fig. 5 and Fig. 11 (Appendix C.2). The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks. This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies (Bansal and Tomlin, 2021; Zhang et al., 2024b). The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs.
@@
PINNs often use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs. The PINN residuals, including PDEs, are further optimized using gradient descent. Smooth activation functions like Tanh support stable and accurate gradient flow throughout training, enabling more effective value approximations. In contrast, VIP with ReLU often encounters zero-gradient regions during backpropagation, which results in gradient explosion for deeper network architectures or degrades the learning of value functions due to insufficient nonlinearity. Therefore, the choice of a smooth activation function like Tanh is better suited for physics-informed learning, thereby ensuring more accurate approximations of the value functions. We also investigate the impact of weight parameters for PINN loss terms during VIP training. Specifically, we evaluate three configurations in Eq. 17: 1) balanced weights for all loss terms; 2) a large weight for anchor loss while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms.
@@
As shown in Fig. 7, the best accumulated reward performance is achieved when all loss terms are properly balanced. Imbalanced weight settings yield stiffness in the training dynamics of PINNs (Wang et al., 2021), which makes the dominant loss term (with the largest weights) converge faster than the others (Wang et al., 2022). In our experiments, such an imbalance causes VIP only to satisfy the PDEs residuals or anchor loss during training, ultimately leading to poor value approximations.
Time discretization impact. Lastly, we evaluate the robustness of VIP and a well-trained MADDPG model (Lowe et al., 2017a) by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts. Notably, in the original discrete-time training setting with a 4-dimensional state and 1-dimensional action space, MADDPG can already achieve near-optimal performance. Fig. 8 illustrates that VIP maintains a nearly constant return across different time intervals, whereas MADDPG’s performance degrades significantly as the interval increases. This result highlights the advantage of VIP in continuous-time multi-agent scenarios.
Figure 9: Comparison between VIP and DTRL baselines.
DTRL algorithms’ comparison. Figure 9 compares VIP with three widely used discrete-time MARL baselines (MATD3, MADDPG, MAPPO) across several multi-agent MuJoCo tasks. We observe that all discrete-time methods’ performance degrades substantially in tasks such as Ant and HalfCheetah. This degradation is not surprising. Discrete-time Bellman backups implicitly assume fixed and sufficiently small time steps, and their value estimates become biased or unstable when the underlying system evolves in continuous-time with arbitrary or environment-dependent time intervals. Moreover, large integration steps amplify approximation errors in both the critic update and the target network, leading to inaccurate temporal-difference objective functions and unstable actor gradients, which aligns well with the observation in other works (Tallec et al., 2019; Park et al., 2021; De Asis and Sutton, 2024).
5 Conclusion
We propose a novel approach that integrates PINNs into the actor-critic framework to solve CT-MARL problems. Specifically, we approximate value functions using HJB-based PINNs and introduce VGI to improve value approximations, thus mitigating the adverse impact of inaccurate value approximations on policy learning. We validate our VIP across continuous‑time variants of MPE and MuJoCo environments and empirically demonstrate that VIP converges faster and achieves higher accumulated reward compared to baselines of SOTA. Furthermore, we investigate the importance of activation function choice and loss term weighting for VIP performance. In summary, our proposed VIP offers a promising approach to solve CT-MARL problems. Since our current formulation is based on the HJB equation, we focus on cooperative settings; extending VIP to HJI-based formulations that handle competitive or mixed-motive scenarios is a compelling direction for future work.
Acknowledgments
This work was supported by the National Science Foundation under Grant No. 2418106.
Ethics Statement
This work focuses on decision-making under continuous-time multi-agent systems. All experiments are conducted entirely in simulation and do not involve human subjects or personal data.